OpenStudy (anonymous):

if f(x)=2x+3, f(0)=? I need some help jogging my memory

OpenStudy (noseboy908):

remember that f(x) simply means "the function "F" of the given "x"

OpenStudy (anonymous):

if you are looking for the f(0) then you need to replace the variable x with a 0 in your initial equation

OpenStudy (noseboy908):

so fill in what you know. the x here is 0. substitute and solve

OpenStudy (anonymous):

so f(0)=3

OpenStudy (noseboy908):

yep

OpenStudy (anonymous):

f(0)=2(0)+3 f(0)=0+3 f(0)=3 y=3

OpenStudy (anonymous):

yeah that's what I got...I'm in college and I can't remember these stupid problems;/ thx

OpenStudy (noseboy908):

welcome.